40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 11 04:32:53 2010, yuta, Summary, CDS, found poor contact of DAC cable, previous A2L results were wrong 
    Reply  Thu Nov 11 11:51:30 2010, Koji, Summary, CDS, found poor contact of DAC cable, previous A2L results were wrong 
       Reply  Thu Nov 11 23:35:23 2010, Koji, Summary, CDS, found poor contact of DAC cable, previous A2L results were wrong IMG_3698.jpgbad-boys.pngIMG_3705.jpgIMG_3706.jpg
Message ID: 3895     Entry time: Thu Nov 11 11:51:30 2010     In reply to: 3891     Reply to this: 3901
Author: Koji 
Type: Summary 
Category: CDS 
Subject: found poor contact of DAC cable, previous A2L results were wrong 

The cause is apparent! The connectors on the cables are wrong!
Currently only 50% of the pin length goes into the connector!

Quote:

(Koji, Jenne, Yuta)

We found one of DAC cables had a poor contact.
That probably caused our too much "tilt" of the beam into MC.

  It was because one of the DAC cables(labeled CAB_1Y4_88) had a poor contact.
  If I push it really hard, it is ok. But maybe we'd better replace the cable.

What caused a poor connection?:
  I don't know.
  A month ago, we checked that they are connected, but things change.

 

ELOG V3.1.3-